### v1

# Contrast-Aware Calibration for Fine-Tuned CLIP: Leveraging Image-Text Alignment

Abstract
Vision-language models (VLMs), such as CLIP, have demonstrated exceptional generalization capabilities and can quickly adapt to downstream tasks through prompt fine-tuning. Unfortunately, in classification tasks involving non-training classes, known as open-vocabulary setting, fine-tuned VLMs often overfit to train classes, resulting in a misalignment between confidence scores and actual accuracy on unseen classes, which significantly undermines their reliability in real-world deployments. Existing confidence calibration methods typically require training parameters or analyzing features from the training dataset, restricting their ability to generalize unseen classes without corresponding train data. Moreover, VLM-specific calibration methods rely solely on text features from train classes as calibration indicators, which inherently limits their ability to calibrate train classes. To address these challenges, we propose an effective multimodal calibration method Contrast-Aware Calibration (CAC). Building on the original CLIP’s zero-shot adaptability and the conclusion from empirical analysis that poor intra-class and inter-class discriminative ability on unseen classes is the root cause, we calculate calibration weights based on the contrastive difference between the original and fine-tuned CLIP. This method not only adapts to calibrating unseen classes but also overcomes the limitations of previous VLM calibration methods that could not calibrate train classes. In experiments involving 11 datasets with 5 fine-tuning methods, CAC consistently achieved the best calibration effect on both train and unseen classes without sacrificing accuracy and inference speed.
Machine Learning, ICML
1 Introduction
Vision-language models, such as CLIP (Radford et al., 2021), pre-trained on vast web-scale text-image datasets, have demonstrated impressive zero-shot capabilities and image-text alignment in various downstream image classification tasks (Deng et al., 2009; Helber et al., 2019). Various prompt learning methods have been proposed to enhance VLM performance on specific tasks using limited labeled data (Zhou et al., 2022b; Khattak et al., 2023a, b). Given CLIP’s strong zero-shot adaptability, the open-vocabulary setting has become a standard for evaluating the performance of fine-tuned VLMs, where prompts are trained on a subset of classes and evaluated on both train and unseen classes (Lee et al., 2023; Tan et al., 2024).
Unfortunately, fine-tuned VLMs regularly overfit to train classes, forgetting the well-calibrated predictions and image-text alignment achieved during pre-training (Zhou et al., 2022b, a). For unseen classes, they often produce semantically unbalanced representations, leading to image-text misalignment and a significant discrepancy between confidence scores and actual accuracy (Guo et al., 2017; Minderer et al., 2021). Existing calibration methods (Joy et al., 2023; Oh et al., 2023; Zadrozny & Elkan, 2001) typically rely on training or analyzing features from the training dataset, limiting their ability to calibrate classes outside the training dataset. Additionally, due to relying on train class text features and overlooking the critical characteristic of image-text alignment, the performance of state-of-the-art (SOTA) VLM’s calibration method, Distance-Aware Calibration (Wang et al., 2024), fails to calibrate the train classes and struggles to handle methods with well-aligned features.
To fundamentally address the miscalibration in fine-tuned CLIP, we conduct extensive empirical analysis, identifying poor intra-class and inter-class discriminative ability on unseen classes caused by downstream task adaptation as the primary cause. Moreover, our experiments reveal that original CLIP trained on large-scale datasets tends to exhibit superior confidence calibration performance, consistent with the findings in (Minderer et al., 2021; Tu et al., 2023). The above findings and the connection between the contrast metric and confidence calibration inspire us to develop Contrast-Aware Calibration (CAC) to achieve effective confidence calibration. CAC improves the intra-class and inter-class discriminative ability of fine-tuned CLIP by reweighting the logits using the well-aligned information from the original CLIP. Specifically, by leveraging the similarity between the logits generated by the original and fine-tuned CLIP, CAC realigns the image-text feature relationships of the fine-tuned CLIP on both train and unseen classes, improving confidence calibration and overcoming the limitations of previous methods that were restricted to train or unseen classes. Notably, as a post-hoc calibration technique, CAC does not affect the model’s accuracy and inference speed, while delivering more reliable predictions.
Due to its design tailored for CLIP with strong image-text alignment, CAC can be applied to any set of classes and any fine-tuning method of CLIP. To assess its confidence calibration effectiveness, we evaluated CAC on both train and unseen classes across 11 datasets, applying it to 5 different prompt learning methods. Leveraging the image-text alignment of the original CLIP, CAC consistently achieves the best performance in confidence calibration, both for train and unseen classes, outperforming the current best calibration method MIR (Roelofs et al., 2022) and SOTA VLM-specific calibration method DAC (Wang et al., 2024). Additionally, CAC consistently enhances the calibration performance of various fine-tuning methods, even for well-aligned methods such as PromptSRC (Khattak et al., 2023b).
In summary, the main contributions of this paper include:
  • To investigate the issue of miscalibration in fine-tuned CLIP on unseen classes, we identify poor intra-class and inter-class discriminative ability as the root cause through empirical analysis, providing a reliable reference for future research.
  • Benefiting from the inference mechanism of contrastive learning models, we establish a connection between the contrast metric and confidence calibration, enabling us to utilize image-text alignment techniques to address the incomplete calibration of previous single-modal calibration methods.
  • Based on the analysis, we propose a novel Contrast-Aware Calibration, which addresses the limitations of previous methods, such as suboptimal calibration and the inability to handle entire classes.
  • The experimental results on 11 datasets, testing the calibration of 5 fine-tuning methods, show that CAC outperforms existing calibration methods. In particular, even for the well-calibrated method, PromptSRC, CAC consistently brings significant improvement.
2 Related Works
Prompt Learning in Vision Language Models. Due to the large parameter size of VLMs and the limited availability of training data for downstream tasks, it is impractical to fine-tune all parameters of the VLMs to adapt them to these tasks. Inspired by the success of prompt learning in NLP (He et al., 2022; Li & Liang, 2021; Liu et al., 2021), many researchers have proposed to adapt VLMs by learning the prompts in end-to-end training. As the pioneering work, CoOp (Zhou et al., 2022b) for the first time introduces the learnable prompt to transfer the task-specific knowledge to VLMs. To improve the generalization of the learnable language prompt in CoOp, CoCoOp (Zhou et al., 2022a) and VPT (Jia et al., 2022) generate a vision-conditional prompt by fusing the image feature and the learnable language prompts. KgCoOp (Yao et al., 2023), ProGrad (Zhu et al., 2023), and other prompt-based methods (Lee et al., 2023; Tan et al., 2024; Yu et al., 2023) are another prompt-based methods for VLMs. MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b) conduct the visual-textual prompt learning by jointly conducting the prompt learning on the vision and text encoders. These fine-tuning methods only train prompts, leading to higher confidence bias for unseen classes. To address this issue, the proposed method leverages the well-calibrated confidence properties of the original CLIP to correct the confidence of fine-tuned VLMs, thus enhancing the reliability of their outputs.
Confidence Calibration. Confidence calibration has been widely studied to ensure that the confidence levels output by models accurately reflect their true performance. A common approach to achieving this is by applying calibration techniques after the model has been trained. These calibration methods can typically be categorized into two main paradigms: scaling-based methods (Guo et al., 2017; Tomani et al., 2022; Yu et al., 2022; Xiong et al., 2023) and bin-based methods (Zadrozny & Elkan, 2001; Zhang et al., 2020; Zadrozny & Elkan, 2002; Roelofs et al., 2022). For scaling-based calibration, a representative method is temperature scaling (Tomani et al., 2022), which learns a single temperature value for rescaling the logits. ATS (Joy et al., 2023) modifies the confidence predicted by per-data-point adaptive temperature. With the wide use of VLMs, several works have explored calibration in VLMs (LeVine et al., 2023; Oh et al., 2023) recently. Distance-Aware Calibration (Wang et al., 2024) estimates the scaling weights for unseen class logits based on textual features, focusing solely on changes in the textual modality. As a result, its ability to handle confidence calibration is limited when encountering fine-tuned VLMs with well-aligned image-text features. In this paper, we propose a novel confidence calibration method, leveraging the image-text alignment of CLIP.
3 Preliminary
Our method is primarily built upon CLIP and its prompt learning methods. Therefore, before introducing the proposed calibration method, we first review the necessary knowledge, including the core concepts of CLIP, prompt learning, and the Expected Calibration Error (ECE) metric.
CLIP. CLIP is developed to align visual and textual data in a common embedding space. CLIP consists of two encoders: an image encoder denoted as $f$ and a text encoder denoted as $g$. During the training phase, the encoders extract feature representations $f(I)$ and $g(E_{w}(T))$ from an input image $I$ and its corresponding text caption $T$, respectively. The term $E_{w}$ represents the word embedding layer, tasked with transforming words into vector representations.
During the zero-shot classification phase, CLIP begins with an image $I$ and a set of hand-designed text captions $[T_{1},T_{2},\dots,T_{N}]$, formatted as $a\ photo\ of\ a\ [CLASS_{i}]$, where $a\ photo\ of\ a$ is a hand-designed template and $[CLASS_{i}]$ specifies a class from $N$ candidate image categories. The image and captions are processed by their respective encoders to extract features, allowing for the computation of class prediction probabilities as follows:
$p(y=i|I)=\frac{\exp(cos(f(I),g(E_{w}(T_{i})))/\tau)}{\sum_{j=1}^{N}\exp(cos(f(I),g(E_{w}(T_{j})))/\tau)}$ (1)
In this context, $\tau$ is the temperature coefficient, and $cos(\cdot,\cdot)$ represents the cosine similarity between features.
Prompt Learning. To effectively adapt VLMs to downstream tasks, prompt learning methods aim to generate more adaptive classifiers, without the need to fine-tune the text encoder $g$. For instance, some studies (Zhou et al., 2022b, a) employ learnable prompts $P=[t_{1},t_{2},\dots,t_{b}]$ to replace hand-designed language prompt templates, where $t$ represents the prompt vector, and $b$ specifies the prompt’s length. Let $[c_{1},c_{2},\dots,c_{N}]$ represent the word embeddings of class names. The corresponding prediction probability is calculated as follows:
$p(y=i|I)=\frac{\exp(cos(f(I),g([P,c_{i}]))/\tau)}{\sum_{j=1}^{N}\exp(cos(f(I),g([P,c_{j}]))/\tau)}$ (2)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the operation of concatenation. For each downstream task, the learnable prompts $P$ are optimized via cross-entropy classification loss during the few-shot learning phase. As a result, updating the language prompt $P$ will adjust the decision boundaries accordingly, utilizing the generated classifier for the downstream tasks.
Expected Calibration Error. To estimate the expected accuracy from finite samples, we group predictions into $M$ interval bins (each of size $1/M$ ) and calculate the accuracy of each bin. Let $B_{m}$ denote the set of indices corresponding to samples whose prediction confidence lies within the specified interval $I_{m}=(\frac{m-1}{M},\frac{m}{M}]$. The accuracy of $B_{m}$ is
$acc(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y_{i}}=y_{i}),$
where $\hat{y_{i}}$ and $y_{i}$ are the predicted and ground-truth class labels for sample $i$. Then we define the average confidence for the bin $B_{m}$ as
$conf(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p_{i}},$
where $\hat{p_{i}}$ is the confidence for sample $i$, which can be calculated by $\hat{p_{i}}=max(p(y=i|I))$. Formally, a perfectly calibrated model satisfies $acc(B_{m})=conf(B_{m})$ for all $m\in\{1,\dots,M\}$. Then ECE is defined as the difference between the accuracy and confidence of all bins, which can be calculated as:
$ECE=\sum_{m=1}^{M}\frac{|B_{m}|}{M}|acc(B_{m})-conf(B_{m})|,$
4 Analysis
In this section, we introduce the contrast metric and its calculation method, exploring the causes of poor confidence calibration in fine-tuned VLMs by examining feature representations of unseen classes. We then analyze the relationship between output logits and the contrast metric in contrastive learning models, which serves as the underlying logic for our method.
4.1 Contrast Metric
Contrast is an indicator used to measure a model’s ability to distinguish between positive and negative samples, which is widely used in contrastive learning (Le-Khac et al., 2020; Ko et al., 2022). Models with strong discriminative capabilities typically exhibit higher contrast scores and vice versa. Specifically, given a similarity matrix $S\in\mathbb{R}^{N\times C}$, where $N$ represents the number of samples, $C$ denotes the number of classes, and $S[i,j]$ indicates the similarity between sample $i$ and class $j$, the contrast metric is computed through the following three components:
[…]
  • Difference Calculation. The contrast metric is calculated as the difference between the average of positive and negative similarities:
    $\text{Contrast}=\frac{1}{N}\sum_{i=1}^{N}s^{+}_{i}-\frac{1}{N}\sum_{i=1}^{N}s^{-}_{i}$ (3)
For VLMs, contrast represents the model’s ability to distinguish between ground-truth and other classes, serving as an indicator of the model’s image-text alignment performance on the current dataset.
4.2 Empirical Study
Although prompt learning methods (Zhou et al., 2022a; Yao et al., 2023) freeze the parameters of the original CLIP, their learnable prompts often cause overfitting on train classes. For instance, KgCoOp (Yao et al., 2023) learns prompts tailored to train classes and ignores vision prompts, resulting in an imbalance between the text representations of unseen classes and the visual representations of the original CLIP, significantly diminishing its ability to accurately represent unseen classes and producing biased contrast scores (Khattak et al., 2023a, b; Wang et al., 2024). We hypothesize that when the representations of fine-tuned VLMs deviate from the pre-trained image-text alignment, their class scores often become biased toward certain categories or exhibit similar scores across multiple classes, losing the pre-trained ability to discriminative intra-class and inter-class samples and causing miscalibration. To investigate this further, we conducted experiments that confirmed our hypothesis, revealing that high inter-class similarity and high intra-class variation correspond to the above two scenarios, respectively, with representative datasets such as FGVCAircraft (Maji et al., 2023) and Food101 (Bossard et al., 2014):
  • Observation 1: Overconfidence caused by inter-class similarity. For FGVCAircraft, where inter-class similarity is high and class boundaries are difficult to distinguish, fine-tuned VLMs tend to misclassify images into a few dominant classes, resulting in overconfident but incorrect predictions, as shown in Figure 1(a). The contrast score results from untruth classes being more similar than ground-truth class, yielding a negative value. Such overconfidence is particularly evident in fine-grained classification tasks with overlapping inter-class features.
  • Observation 2: Underconfidence caused by intra-class variation. For Food101, where inter-class similarity is low but significant intra-class variation exists, fine-tuned VLMs often produce multiple high-confidence predictions across unseen classes, leading to lower overall confidence, as shown in Figure 1(b). Despite this, the model can still identify the correct category, keeping the contrast score positive.
In summary, we found that low contrast often indicates overconfidence, while increasing contrast transitions the model from overconfidence to underconfidence, eventually achieving proper calibration. We conducted similar experiments on various fine-tuning methods, all of which yielded extremely similar results. As representative single-modal and cross-modal fine-tuning methods, the results of KgCoOp and MaPLe are shown in Figure 1(c) and Figure 1(d). Overall, higher contrast scores are generally linked to lower ECE values, although the relationship between these two metrics can vary between different datasets. Through extensive experimental analysis, we conclude that contrast and ECE exhibit a negative correlation for unseen classes, making contrast a reliable metric for scaling ECE. Furthermore, the conclusions indicate that well-aligned VLMs typically exhibit better confidence calibration, consistent with the observations in (Minderer et al., 2021; Tu et al., 2023).
(a) ECE of FGVCAircraft
(b) ECE of Food101
(c) KgCoOp
(d) MaPLe
Figure 1: (a) & (b) The reliability of KgCoOp evaluated on the FGVCAircraft and Food101 datasets. (c) & (d) The relationship between contrast and ECE for the logits output by KgCoOp and MaPLe across 11 datasets.
4.3 Logits and Contrast
For VLMs like CLIP, which are trained and make predictions based on contrastive learning, the similarity matrix $S$ used to calculate the contrast corresponds to the logits output by CLIP for each sample:
$S=logits_{CLIP}=\frac{f(I)\cdot g(E_{w}(T))}{\|f(I)\|\|g(E_{w}(T))\|}$ (4)
where these symbols are consistent with the definitions provided in the preliminary section. When the contrast is low, the logits from CLIP often indicate two or more classes with relatively high scores. Conversely, a higher contrast typically corresponds to a class with a significantly higher score, while the scores for other classes remain much lower. This property is generally applicable to other CLIP-based models and helps establish the relationship between the logits output by VLMs and the contrast metric. By linking logits with the contrast metric, we transform our conclusions into a practical solution for contrastive learning models, providing the foundational logic for our method.
5 Contrast-Aware Calibration
In this section, we first introduce a unique scaling weight computation method, Contrast-Aware Weights (CAW), specifically designed for image-text alignment models, based on the conclusions from Section 4.2. Then, we propose a more refined calibration method, Contrast-Aware Calibration (CAC), which serves as our method. Finally, we provide a brief overview of CAC calibration process during inference and analyze its advantages.
5.1 Contrast-Aware Weights
As demonstrated in Section 4.2 and existing studies (Minderer et al., 2021; Tu et al., 2023), the original CLIP, trained on large-scale image-text pairs, exhibits relatively high contrast across different datasets, leading to more conservative predictions compared to fine-tuned CLIP. This observation motivates us to leverage original CLIP as a reference to regulate the confident bias logits of fine-tuned CLIP. Specifically, we compute the $L_{1}$ distance between the logits of the original CLIP and fine-tuned CLIP, treating it as the confidence bias metric $z$. The formulaic expression is as follows:
$z=\frac{1}{N}\sum_{i=1}^{N}|P_{i}-\hat{P_{i}}|,$ (5)
where $N$ denotes the total number of classes, and $\hat{P}=\{\hat{p}_{i}\}_{i=1}^{N}$ and $P=\{p_{i}\}_{i=1}^{N}$ represent the logits output by the original and fine-tuned CLIP, respectively. In particular, according to our analysis, the logits of contrastive learning-based VLMs are equivalent to the contrast metric, so $z$ can serve as an indicator for measuring the confidence difference between the original and fine-tuned VLMs. To better leverage its negative correlation with ECE, we design the following function to transform $z$ into CAW:
$\gamma=\alpha\cdot e^{-kz}.$ (6)
This function addresses several issues caused by directly using $z$ as a confidence weight, such as reversed monotonic trends, minimal numerical differences, and an incomplete value range. The components of the function serve the following purposes:
  • Reason for choosing $e^{-x}$: Due to the negative correlation between ECE and contrast, we choose a decreasing function $e^{-x}$, which has a range of $[0,1]$ and aligns with the required monotonicity.
  • Effect of $k$: Since the text and image features in CLIP-based models undergo normalization before computing the logits, the $L_{1}$ distance may be small. Therefore, it is necessary to amplify the input to the function, with $k$ serving to scale $z$, and enable the function to capture input variations more effectively.
  • Effect of $\alpha$: Since fine-tuned VLMs may be underconfidence and overconfident in various datasets, we modify the function’s maximum value to $\alpha~{}(>1)$, equipping CAW with the ability to deal with underconfidence.
In summary, the design of Equation 5 bridges the gap between contrast and the degree of miscalibration, providing a reliable metric to assess the CLIP’s confidence calibration. By applying amplification and other operations, Equation 6 transforms $z$ into a suitable scaling factor, ultimately resulting in the CAW.
5.2 Contrast-Aware Calibration
The CAW proposed in the previous section can already be used as an effective confidence calibration weight for fine-tuned CLIP. However, different datasets and fine-tuning methods require varying levels of calibration. For example, KgCoOp requires stronger calibration, whereas PromptSRC typically only needs minor adjustments. Therefore, we use a piecewise function to amplify the weights in underconfident cases and reduce them in overconfident cases, ultimately forming the CAC method:
$\hat{\gamma}=\begin{cases}\gamma^{2},&\text{if }\gamma<\lambda_{1},\\ \gamma,&\text{if }\lambda_{1}\leq\gamma\leq\lambda_{2},\\ \gamma^{2}&\text{if }\gamma>\lambda_{2}.\end{cases}$ (7)
where $\hat{\gamma}$ represents the final calibration weight, $\gamma$ represents the output of CAW, and $\lambda_{1}$ and $\lambda_{2}$ represent the boundary points of the interval for shrinking or amplifying $\gamma$. Through the following two designed modules, CAC achieves more flexible confidence calibration compared to CAW:
  • Advantage of Thresholds: When $\gamma$ is less than or greater than these thresholds, the model often exhibits significant underconfidence or overconfidence, respectively. Therefore, we select manually defined thresholds to regulate scaling weights in specific scenarios.
  • Advantage of Squaring: For values within the range $[0,\alpha]$, squaring will offer the advantage of maintaining numerical stability without drastically affecting the values. Unlike scaling by a manually defined constant, this approach removes the need for hyperparameter fine-tuning, making it more efficient and reliable.
Ultimately, the proposed CAC is a robust confidence calibration method specifically tailored for CLIP with well-aligned image-text features. By scaling the logits of fine-tuned CLIP using those of the original CLIP, CAC achieves parameter efficiency, strong generalization capabilities, and a training-free implementation, addressing the limitations of previous methods that rely on training parameters or analyzing train data features.
5.3 Calibrated inference
Given an input image $i$, we first collect the CAC scores of this image, denoted as $\hat{\gamma}_{i}$, which is then used to calculate the rectified logits as follows:
$L^{CAC}_{i}=\hat{\gamma}_{i}*\tau*logits_{i}$ (8)
where $logits_{i}$ is the logit calculated by fine-tuned CLIP. As a post-hoc confidence calibration method, CAC exclusively scales the temperature coefficient $\tau$ of CLIP, improving the model’s reliability without compromising its original accuracy and inference speed. Notably, leveraging the open-vocabulary classification capabilities of the original CLIP, CAC can automatically adapt to any input class and fine-tuning method, enabling automatic calibration for both train and unseen classes in various fine-tuning methods.
In summary, the advantages of our method are as follows:
  • Tailored for CLIP: CAC is specifically designed for CLIP, leveraging the original CLIP’s alignment between visual and textual modalities to achieve effective confidence calibration.
  • Simultaneous Scaling for Both Train and Unseen Classes: As an improvement over previous methods, CAC calibrates the confidence of both train and unseen classes, addressing their limitation of only handling train or unseen classes.
  • Strong Empirical Foundation: The design of CAC is grounded in experimental insights, ensuring its strong interpretability and reliability.
  • Simple Plug-and-Play: CAC eliminates the need for additional training or extensive analysis of train data, providing a straightforward and efficient solution for enhancing confidence calibration in VLMs.
6 Experiments
6.1 Experimental Setup
Evaluation Paradigm. Following the open-vocabulary setting in the VLM field (Zhou et al., 2022a; Khattak et al., 2023a), the datasets are divided into train and unseen classes. The model is trained on the train classes in a few-shot setting and we generally report the calibration performance on both train and unseen classes, which is different from DAC (Wang et al., 2024).
Compared Methods. We mainly focus on benchmarking against the other 5 current representative prompt learning methods: CoCoOp (Zhou et al., 2022a), KgCoOp (Yao et al., 2023), MaPLe (Khattak et al., 2023a), ProGrad (Zhu et al., 2023), and PromptSRC (Khattak et al., 2023b). Since models like CoOp only consider the text modality and exhibit low accuracy, making their calibration significance minimal, we focused on testing and calibrating its optimized version, CoCoOp and KgCoOp. For train classes calibration, we select three representative calibration methods: Histogram Binning (HB) (Zadrozny & Elkan, 2001), Isotonic Regression (IR) (Zadrozny & Elkan, 2002), and Multi-Isotonic Regression (MIR) (Roelofs et al., 2022). For unseen classes calibration, we compare the SOTA method designed for CLIP, DAC (Wang et al., 2024). Datasets and Implementation details used in the experiments are shown in Appendix A.
Implementation details. In the main experiments of this paper, we set $k=15$ and $\alpha=1.10$ as the default parameters and select $\lambda_{1}=0.9$ and $\lambda_{2}=1.0$ as the threshold values. Additional details about pre-trained models, hyperparameters, and implementation specifics are provided in Appendix A.
Evaluation metrics. We used 4 standard metrics in our evaluation of open-vocabulary confidence calibration: Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019) and Proximity Informed Expected Calibration Error (Xiong et al., 2023). All of the calibration error is given by ×$\times 10^{-2}$.
6.2 Main Results
How effective is CAC calibration of VLMs output for unseen classes? As shown in Table 1, CAC achieves optimal performance across all datasets, highlighting the advantages of image-text alignment calibration. Notably, CAC performs well on prompt learning methods such as KgCoOp and ProGrad, showing significant performance improvements, whereas DAC demonstrates the opposite trend in performance optimization. Moreover, for methods like PromptSRC, which already surpass the original CLIP in calibration performance, CAC further reduces its ECE from 4.29 to 3.47, demonstrating the strong calibration capability inherent in the logits of the original CLIP. In summary, the comparison between DAC and CAC across multiple methods and four metrics validates the effectiveness of image-text alignment supported by an empirical foundation.
Table 1: Average calibration performance across 11 datasets. “Conf” represents the original performance on open-vocabulary classes with existing tuning methods. $\downarrow$ indicates smaller values are better. Bold numbers are significantly superior results.
Method ECE($\downarrow$) ACE($\downarrow$) MCE($\downarrow$) PIECE($\downarrow$)
Conf DAC CAC Conf DAC CAC Conf DAC CAC Conf DAC CAC
CoCoOp 5.44 5.70 4.24 5.35 5.60 4.22 1.38 1.40 1.20 7.35 8.06 6.83
KgCoOp 3.98 4.11 3.85 3.93 4.09 3.78 1.08 1.18 1.10 6.45 6.62 6.39
MaPLe 7.80 5.91 5.35 7.77 5.93 5.30 2.08 1.62 1.61 9.53 8.19 7.69
ProGrad 5.04 6.13 4.04 4.95 6.18 4.05 1.47 1.53 1.24 7.41 8.20 6.75
PromptSRC 4.29 4.55 3.47 4.24 4.41 3.40 1.16 1.17 1.03 6.70 6.82 6.12
How effective is CAC calibration of VLMs output for train classes? As shown in Table 2, CAC surpasses existing calibration methods and achieves SOTA calibration performance in the train classes. Notably, some traditional calibration methods perform poorly in the VLM field, with ECE increasing instead of decreasing, highlighting the importance of designing calibration methods specifically tailored for VLMs. Compared to the current SOTA method MIR (Roelofs et al., 2022), CAC demonstrates significantly better performance across all datasets, even on the well-calibrated fine-tuning method PromptSRC, with improvements ranging from 0.19 to 0.99. In summary, due to its design based on the well-calibrated original CLIP, CAC has the capability to calibrate any class in open-vocabulary settings, ensuring robust and reliable calibration. We provide detailed experimental results in Appendix C.
Table 2: Average calibration performance across 11 datasets on train classes. “Conf” represents the origin performance without existing calibration methods.
Method Conf IR HB MIR CAC
CoCoOp 3.60 7.80 7.50 3.87 3.05
KgCoOp 5.87 7.23 7.41 7.38 4.10
 MaPLe 2.80 7.81 6.63 2.70 2.50
ProGrad 5.93 5.69 5.63 4.42 4.17
PromptSRC 3.74 6.39 6.38 3.55 2.75
6.3 Ablation Studies
In this section, we first conducted ablation experiments to explore the effect of each module in CAC. We then provided a detailed analysis of the underlying principles behind the optimal parameter choices for each module. Finally, we compared and analyzed the effects of the piecewise function parameters, confirming that CAW is the key factor responsible for achieving good confidence calibration.
The Effect of Each Component in CAC. We evaluate the effect of each component in CAC on calibration performance by systematically ablating the designed modules and assessing their performance on test datasets as follows:
  • Without $\alpha$: As analyzed earlier, when $\alpha$ is removed, CAC struggles to handle underconfident datasets, resulting in poor performance. Therefore, the inclusion of $\alpha$, which expands the range of values, is essential.
  • Without $k$: As shown in Table 4, ablating $k$ hinders the model’s ability to effectively distinguish the contrast differences, resulting in worse calibration, which aligns with our initial design rationale. However, omitting normalization when calculating logits can result in a numerical explosion, further exacerbating miscalibration. Therefore, using amplification based on $k$ remains the optimal choice.
  • Without the EXP function: We conducted comparative experiments under the same settings by replacing the EXP function with the monotonic decreasing $1/x$ function. Since the value of $1/x$ approaches infinity as $x$ approaches zero, it fails to handle well-aligned methods like PromptSRC, thereby demonstrating the rationale behind using the EXP function.
  • Without the piecewise function: Since the piecewise function is primarily designed for datasets requiring significant confidence calibration, its ablation has a less pronounced effect on overall performance compared to other modules. The results indicate that most datasets benefit from stronger scaling, validating the necessity of the piecewise function in our design.
Table 3: Average calibration performance of different $k$ of CAC across 11 datasets.
Method Conf 10 15 20 25
CoCoOp 5.44 4.82 4.24 6.41 11.09
KgCoOp 3.98 3.82 3.85 3.66 5.16
MaPLe 7.80 6.82 5.35 8.14 12.52
ProGrad 5.04 4.57 4.04 8.20 12.74
PromptSRC 4.29 3.93 3.47 5.33 8.31
Table 4: Average calibration performance across 11 datasets in 5 prompt learning methods. “w/o” indicates the results of CAC without the inclusion of that specific component.
Method Conf DAC w/o $k$ w/o $\alpha$ w/o EXP w/o piecewise CAC
CoCoOp 5.44 5.70 9.08 9.73 11.15 4.72 4.24
KgCoOp 3.98 4.11 7.52 8.12 26.74 3.97 3.85
MaPLe 7.80 5.91 11.32 11.25 28.13 6.57 5.35
ProGrad 5.04 6.13 8.92 11.84 28.85 4.44 4.04
PromptSRC 4.29 4.55 7.80 9.41 23.87 3.84 3.47
The Impact of Different $k$. Intuitively, scaling the CAW value to an appropriate range produces optimal confidence calibration results, while excessively high or low values lead to suboptimal performance. Table 3 confirms our hypothesis, showing that a proper value of $k$ achieves the best calibration performance. In general, this experiment highlights that the regularized $L_{1}$ distance tends to be relatively small, requiring appropriate amplification for optimal results.
The impact of different $\alpha$ values. As shown in Table 5, we tested the effect of 4 different $\alpha$ values, with the default setting of $\alpha$ being $1.10$. When $\alpha=1.00$, the model cannot handle underconfidence, resulting in degraded calibration performance. As $\alpha$ increases, CAC’s ability to mitigate underconfidence improves. However, when $\alpha\geq 1.20$, the cases of underconfidence are largely neglected, leading to CAC calibration to change to overconfidence, which compromises its balance between underconfidence and overconfidence and leads to suboptimal results. In summary, as intended in our design, $\alpha$ plays a critical role in the model’s ability to address both underconfidence and overconfidence.
Table 5: Average calibration performance of different $\alpha$ of CAC across 11 datasets.
Method Conf 1.00 1.05 1.10 1.20
CoCoOp 5.44 7.87 5.88 4.24 4.84
KgCoOp 3.98 5.86 3.81 3.85 4.73
MaPLe 7.80 9.13 6.93 5.35 5.85
ProGrad 5.04 10.38 7.09 4.04 4.29
PromptSRC 4.29 7.87 5.00 3.47 3.90
Impact of Piecewise Function Thresholds on Confidence Calibration. For the overconfidence threshold $\lambda_{1}$, we set 0.90 as the default scaling coefficient. As shown in Table 6, values below $\lambda_{1}$ indicate a significant deviation between the model output and the original CLIP. For the underconfidence threshold $\lambda_{2}$, we hypothesize that when the scores of a fine-tuned CLIP closely align with those of the original CLIP, its predictions are more reliable; therefore, we select 1.00 as the underconfidence threshold. We also tested two alternative thresholds, 0.95 and 1.05, which showed minimal differences in results, suggesting that most outputs resembling the original CLIP’s predictions are underconfident. In summary, the minimal impact of the piecewise function thresholds highlights that CAW is the key factor in achieving excellent confidence calibration, further validating its effectiveness. We also analyze the effect of model backbone on confidence calibration in Appendix B.
Table 6: The impact of different piecewise function thresholds on CAC confidence calibration.
Method $\lambda_{1}$ $\lambda_{2}$ CAC
0.85 0.95 0.95 1.05
CoCoOp 5.63 4.84 4.63 4.57 4.24
KgCoOp 4.22 3.55 3.72 3.59 3.85
MaPLe 7.88 5.92 5.71 5.69 5.35
ProGrad 5.19 5.27 4.71 4.65 4.04
PromptSRC 4.41 4.12 3.83 3.69 3.47
7 Conclusion
Prompt learning for vision-language models (VLMs) has gained significant attention; however, fine-tuned VLMs face substantial calibration challenges in open-vocabulary settings. Existing calibration methods struggle to address unseen classes, while recently proposed VLM-specific calibration approaches encounter inability to handle train classes and suboptimal performance. In this paper, we present a comprehensive study on VLM calibration in open-vocabulary settings. Through empirical analysis, we identify poor intra-class and inter-class discriminative ability on unseen classes as the primary cause of miscalibration. To address this, we establish a connection between contrast metric and confidence calibration and then propose the Contrast-Aware Calibration (CAC) method. Extensive experimental results demonstrate that our proposal consistently achieves SOTA calibration performance on both train and unseen classes, without compromising accuracy or inference speed. Moreover, CAC improves the calibration effectiveness of various prompt learning methods across the board.
One limitation is the need for manual parameter tuning for each module and we will explore automated parameter tuning methods to provide deeper insights.
Impact Statement
This paper aims to enhance the confidence calibration performance of fine-tuned CLIP, thereby increasing reliability in real-world deployments. While there are many potential societal consequences of our work, we firmly believe that the majority of these impacts are positive, and we do not find it necessary to highlight any specific ones here.
Appendix A Experimental Setting
Dataset. Following the CoCoOp framework (Zhou et al., 2022a), we conducted evaluations of our proposed DAC along with comparison calibration methods on various image classification tasks. These tasks included general object recognition using ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) datasets, fine-grained object recognition involving datasets such as Oxford Pets  (Parkhi et al., 2012), Food-101 (Bossard et al., 2014), StanfordCars (Krause et al., 2013), Oxford Flowers102 (Nilsback & Zisserman, 2008), and FGVCAircraft (Maji et al., 2023). Additionally, we performed a remote sensing recognition task using the EuroSAT  (Helber et al., 2019) dataset, a texture recognition task using the DTD (Cimpoi et al., 2014) dataset, an action recognition task using UCF101 (Soomro et al., 2012) dataset and a large-scale scene understanding task using SUN397 (Xiao et al., 2010) dataset.
Implementation details. We use CLIP (ViT-B/16) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022b). Following the official implementation, We list the general hyperparameters in Table 7. Here, we briefly introduce the corresponding exclusive hyperparameters of each VLM tuning method. All the methods are adopted from their official implementation. For CoCoOp, they do not contain other hyperparameters. For ProDA, we set $\lambda=0.1$ For KgCoOp, we set $\lambda=8.0$. For MaPLe, we set prompt depth $J$ to 0 and the language and vision prompt lengths to 2. For ProGrad, we set $\lambda=0.8$. For PromptSRC, we set deep prompting with V = T = 4. $\lambda_{1}=10$ and $\lambda_{2}=25$ are used in weight loss. For textual diversity, we use a total of $N=60$ standard prompt templates. We reproduce their methods in our framework based on the standard codebase.
Table 7: Hyper-parameter settings for different models. All of fine-tuning method is rigorously reproduced based on the experimental details given in its paper.
Method Prompt Depth Prompt Length Epochs Warmup Epochs Learning Rate Batch Size
CoCoOp 1 4 20 1 0.002 1
KgCoOp 1 16 200 1 0.002 32
MaPLe 9 2 5 1 0.0026 4
ProGrad 1 16 100 1 0.002 32
PromptSRC 9 4 50 1 0.0025 4
Appendix B The Impact of Different backbones
Follow DAC (Wang et al., 2024) experimental settings, we evaluated the performance of DAC and CAC on fine-tuned VLMs with different architectures using the OxfordFlowers (Nilsback & Zisserman, 2008) dataset in this section. Unlike accuracy, which tends to increase as the number of shots increases (Zhou et al., 2022a), calibration errors are prevalent across diverse few-shot settings. As shown in Table 8, CAC demonstrates stronger cross-model generalization capabilities, achieving effective calibration across all architectures.
Table 8: Comparison results of ECE (%) using different visual backbones on OxfordFlowers dataset. The smaller values are better.
Method RN50 ViT-B/32 ViT-B/16
Conf DAC CAC Conf DAC CAC Conf DAC CAC
CoCoOp 6.62 4.74 3.87 9.71 6.08 4.11 7.16 5.94 4.21
KgCoOp 3.27 4.38 4.25 4.94 5.06 4.88 4.80 4.38 5.10
MaPLe 4.14 3.20 3.34 8.95 5.33 4.8 14.67 8.60 5.57
ProGrad 16.23 3.62 3.54 5.87 4.49 4.44 5.30 4.42 3.77
PromptSRC 3.97 4.28 3.83 5.63 5.10 5.00 5.92 5.59 5.83
Appendix C Detailed Experimental Results
In this section, we illustrate the experimental results in detail.
C.1 The calibration performance of fine-tuned CLIP
To showcase the miscalibration in fine-tuned VLMs, we fine-tune the pre-trained CLIP with 5 tuning methods on 11 downstream datasets and use ECE as the calibration metric. We report the detailed calibration results in this section. As is shown in Table 9(b), we can observe that distillation-based tuning (KgCoOp, ProGrad and PromptSRC) restrains the model’s confidence in train classes. For the calibration of the unseen class, distillation-based tuning utilizes zero-shot CLIP as the teacher model to improve the generalization ability, which generates relatively reliable predictions in unseen classes as expected. More importantly, we observe that fined-tuned VLMs tend to be underconfident in train classes while prone to overconfident in unseen classes. This phenomenon motivates us to investigate the calibration of fine-tuned VLMs further.
Table 9: Expected Calibration Error (ECE) on diverse downstream datasets using various tuning methods for CLIP-ViT-B/16. The calibration performance is averaged across three variants. Caltech, OP, SCars, Flowers, Food, FGVC, SUN, and UCF represent Caltech101, OxfordPets, StanfordCars, Flower102, Food101, FGVCAircraft, SUN397, and UCF101.
(a) Base
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 6.48 4.71 8.29 2.71 1.57 1.53 3.09 2.25 3.80 1.54 3.43 3.58
CoCoOp 1.45 2.20 7.19 3.82 0.87 2.65 7.74 2.33 7.68 1.51 2.17 3.60
KgCoOp 2.42 7.45 11.14 6.81 1.38 2.64 9.91 2.93 10.90 4.88 4.16 5.87
MaPLe 1.60 3.47 2.77 4.27 0.73 1.89 4.07 2.16 7.20 0.95 1.72 2.80
ProGrad 2.94 11.86 9.94 6.51 1.41 2.66 9.14 2.97 9.45 3.74 4.56 5.93
PromptSRC 2.37 3.10 9.17 3.25 0.82 2.12 4.75 2.88 8.28 2.08 2.35 3.74
(b) Unseen
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.59 9.45 9.18 6.57 1.79 2.10 4.94 3.42 3.22 3.55 5.28 4.64
CoCoOp 2.44 13.27 10.20 13.43 1.91 1.62 7.16 2.31 2.09 1.73 3.73 5.44
KgCoOp 1.47 5.10 7.18 10.74 1.94 1.91 4.80 3.28 3.22 1.21 2.97 3.98
MaPLe 2.45 20.77 17.18 16.52 1.16 2.99 14.67 2.59 3.01 2.10 2.41 7.80
ProGrad 2.33 5.81 16.81 10.19 1.52 2.55 5.30 2.89 3.46 1.06 3.48 5.04
PromptSRC 1.85 4.83 7.79 15.41 1.53 1.63 5.92 2.98 1.87 0.94 2.47 4.29
C.2 Calibrated performance for fine-tuned CLIP
We utilize commonly used post-hoc confidence calibration for fine-tuned CLIP and expect to calibrate them well on both train and unseen classes. We can observe that established calibration can remedy miscalibration in train classes. However, such efficacy can not transfer to unseen classes.
C.2.1 TRAIN CLASS
Post-hoc calibration can remedy miscalibration in train classes. In the latest works (Zadrozny & Elkan, 2001, 2002), scaling-based methods like TS have been successfully adopted in VLM calibration, and their findings reveal that this calibration can handle miscalibration well in the close-world setting. As a complement, we observe that bin-based (e.g., MIR) calibration can notably reduce miscalibration in train classes. Post-hoc calibration on train classes can not be transferred to unseen classes. For instance, scaling-based methods like TS scale the logits by a single scalar temperature value. If fine-tuned VLMs show underconfidence in train classes, TS sharpens the logit scale before the softmax function and makes the prediction more confident. As a cost, the confidence distribution of fine-tuned VLMs become sharper and causes these models to display exaggerated overconfidence in unseen classes. for bin-based methods, they need the probabilities from train classes as the input, which are incompatible with zero-shot prediction.
C.2.2 UNSEEN CLASS
Table 10: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
(a) W/O Calibration
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.59 9.45 9.18 6.57 1.79 2.10 4.94 3.42 3.22 3.55 5.28 4.64
CoCoOp 2.44 13.27 10.20 13.43 1.91 1.62 7.16 2.31 2.09 1.73 3.73 5.44
KgCoOp 1.47 5.10 7.18 10.74 1.94 1.91 4.80 3.28 3.22 1.21 2.97 3.98
MaPLe 2.45 20.77 17.18 16.52 1.16 2.99 14.67 2.59 3.01 2.10 2.41 7.80
ProGrad 2.33 5.81 16.81 10.19 1.52 2.55 5.30 2.89 3.46 1.06 3.48 5.04
PromptSRC 1.85 4.83 7.79 15.41 1.53 1.63 5.92 2.98 1.87 0.94 2.47 4.29
(b) DAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.59 9.45 9.18 6.57 1.79 2.10 4.94 3.42 3.22 3.55 5.28 4.64
CoCoOp 2.18 6.26 9.30 3.52 3.81 7.85 5.94 3.88 6.42 8.28 5.26 5.70
KgCoOp 1.44 4.03 9.11 9.32 2.55 1.90 4.38 3.55 3.57 1.58 3.77 4.11
MaPLe 1.91 6.25 10.65 8.53 3.38 2.57 8.60 4.23 4.91 6.55 7.46 5.91
ProGrad 2.88 10.67 8.44 3.03 4.24 3.12 4.42 5.28 6.04 8.42 10.93 6.13
PromptSRC 2.10 4.53 9.59 12.48 2.58 1.59 5.59 3.55 2.51 2.81 2.71 4.55
(c) CAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 2.17 8.84 7.08 10.99 0.77 5.92 5.89 1.87 1.35 2.38 2.35 4.51
CoCoOp 2.45 7.80 9.71 8.62 1.74 1.54 4.21 3.48 2.68 1.20 3.18 4.24
KgCoOp 2.02 5.29 6.39 8.96 0.30 4.85 5.10 2.34 2.90 1.14 3.01 3.85
MaPLe 2.26 10.31 16.22 8.63 1.21 2.30 5.57 2.95 4.52 1.96 2.94 5.35
ProGrad 2.28 5.17 10.7 5.6 1.64 2.15 3.77 3.69 3.39 2.03 4.01 4.04
PromptSRC 1.79 4.14 7.82 7.89 0.92 2.22 5.83 2.29 1.86 1.24 2.19 3.47
Table 11: Expected Calibration Error (ECE) on train classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
(a) W/O Calibration
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 6.48 4.71 8.29 2.71 1.57 1.53 3.09 2.25 3.80 1.54 3.43 3.58
CoCoOp 1.45 2.20 7.19 3.82 0.87 2.65 7.74 2.33 7.68 1.51 2.17 3.60
KgCoOp 2.42 7.45 11.14 6.81 1.38 2.64 9.91 2.93 10.90 4.88 4.16 5.87
MaPLe 1.60 3.47 2.77 4.27 0.73 1.89 4.07 2.16 7.20 0.95 1.72 2.80
ProGrad 2.94 11.86 9.94 6.51 1.41 2.66 9.14 2.97 9.45 3.74 4.56 5.93
PromptSRC 2.37 3.10 9.17 3.25 0.82 2.12 4.75 2.88 8.28 2.08 2.35 3.74
(b) Isotonic Regression
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 3.11 7.39 10.41 8.30 6.34 2.04 4.85 6.58 9.20 7.82 11.73 7.07
CoCoOp 4.90 5.79 8.11 15.66 10.55 1.48 6.88 6.68 9.16 7.65 8.92 7.80
KgCoOp 1.29 4.80 11.50 11.93 9.81 0.99 18.49 8.78 3.80 4.57 3.54 7.23
MaPLe 4.00 7.60 6.87 15.29 12.39 1.47 2.73 7.69 8.85 7.95 11.06 7.81
ProGrad 2.90 3.78 6.61 13.41 8.10 1.19 5.71 3.64 4.84 5.39 7.05 5.69
PromptSRC 2.28 3.78 5.98 14.33 10.78 1.41 4.89 4.62 5.12 6.91 10.22 6.39
(c) Histogram Binning
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 2.16 11.40 9.76 2.23 2.87 4.73 8.24 1.61 8.65 5.03 7.81 5.86
CoCoOp 3.18 17.48 7.37 4.04 2.76 4.68 17.03 5.93 5.98 6.21 7.89 7.50
KgCoOp 0.76 13.78 14.73 4.80 2.38 4.37 15.67 5.16 5.52 5.37 9.00 7.41
MaPLe 0.94 12.45 13.01 7.28 2.90 4.28 14.31 1.32 7.25 3.68 5.46 6.63
ProGrad 1.08 11.66 4.20 5.05 2.41 4.06 13.51 1.94 6.46 5.35 6.22 5.63
PromptSRC 1.05 14.46 13.01 4.02 2.37 4.11 16.20 1.60 4.69 3.97 4.70 6.38
(d) Multi-Isotonic Regression
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 3.47 3.68 2.84 1.97 2.80 0.31 2.90 2.56 1.78 1.93 4.45 2.61
CoCoOp 0.38 5.07 4.09 4.80 2.06 1.43 10.33 1.83 6.97 2.42 3.24 3.87
KgCoOp 0.74 12.33 19.93 5.27 1.27 2.11 16.35 3.76 8.30 5.31 5.80 7.38
MaPLe 0.35 4.02 3.12 4.68 2.87 0.98 4.24 1.62 3.82 2.13 1.90 2.70
ProGrad 0.79 9.26 5.19 4.94 1.46 0.93 11.75 2.02 5.41 2.95 3.89 4.42
PromptSRC 0.60 4.14 10.26 4.11 1.57 0.69 7.04 1.14 6.07 1.60 1.88 3.55
(e) CAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 4.93 4.85 9.11 2.76 0.43 2.30 4.69 1.19 1.38 3.57 4.05 3.57
CoCoOp 1.10 3.44 5.83 2.97 0.46 3.29 6.22 1.61 5.61 1.09 1.94 3.05
KgCoOp 1.70 4.79 9.16 4.67 0.50 1.12 7.91 1.94 8.49 2.21 2.60 4.10
MaPLe 1.26 4.78 2.29 3.96 0.64 1.45 3.04 1.53 5.24 1.76 1.58 2.50
ProGrad 2.30 8.81 7.97 4.26 0.46 1.09 7.19 2.15 7.31 1.40 2.89 4.17
PromptSRC 1.69 2.77 7.25 2.94 0.61 0.94 3.66 1.63 6.42 0.71 1.58 2.75
C.3 Accuracy & Confidence of fine-tuned CLIP
In this section, we analyze the miscalibration from the classification performance. Since open-vocabulary classes are not explicitly seen during the tuning, some might believe that this miscalibration arises from rapid accuracy decreasing and leads to a larger Expected Calibration Error (ECE). We report the average performance on 11 datasets in Table 12(b). Although the classification performance on unseen classes may be comparable or lower, the average predictive confidence of fine-tuned VLMs is still relatively higher than zero-shot CLIP, which results in a larger ECE. The results are the same as for DAC, although fine-tuning improves the accuracy of VLM in downstream tasks, the corresponding ECE also increases.
Table 12: Accuracy comparison of existing prompt tuning in the open-vocabulary setting.
(a) Base
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 97.16 53.24 57.00 27.61 90.06 72.42 71.70 91.33 63.77 69.32 70.89 69.50
CoCoOp 97.91 77.16 85.61 35.07 90.53 75.89 94.24 95.04 71.67 79.29 81.94 80.40
KgCoOp 97.89 80.71 89.51 39.36 90.61 75.98 96.20 94.95 75.07 80.95 84.33 82.32
MaPLe 98.00 80.40 89.25 33.29 90.48 76.57 96.39 95.66 72.87 81.01 84.01 81.63
ProGrad 98.39 76.43 90.02 41.16 90.38 76.96 96.17 94.95 78.52 81.18 84.85 82.64
PromptSRC 98.39 83.49 92.96 44.36 90.62 78.00 98.10 95.48 80.65 83.02 87.63 84.79
(b) Unseen
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 94.14 60.87 64.00 35.93 91.15 68.11 77.50 97.15 74.94 75.59 78.42 74.35
CoCoOp 93.52 55.35 63.04 32.81 91.53 70.41 72.17 97.44 73.50 76.61 73.77 72.74
KgCoOp 94.36 51.29 69.22 30.59 91.57 69.58 73.21 97.45 74.12 75.18 74.09 72.79
MaPLe 93.92 46.26 52.13 22.32 91.85 69.67 67.50 97.73 73.49 77.85 78.73 70.13
ProGrad 93.60 52.05 55.10 31.45 89.40 67.06 73.43 97.07 69.16 72.29 70.85 70.13
PromptSRC 94.07 62.88 74.13 27.99 91.37 70.33 77.23 97.18 75.03 78.96 78.83 75.27
Table 13: Confidence comparison of existing prompt tuning in the open-vocabulary setting.
(a) Base
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 0.92 0.56 0.66 0.29 0.90 0.74 0.76 0.90 0.63 0.73 0.75 0.71
CoCoOp 0.97 0.81 0.82 0.35 0.92 0.78 0.90 0.95 0.68 0.82 0.84 0.80
KgCoOp 0.97 0.78 0.82 0.37 0.92 0.78 0.90 0.94 0.69 0.81 0.84 0.80
MaPLe 0.97 0.86 0.90 0.34 0.92 0.79 0.94 0.95 0.69 0.84 0.86 0.82
ProGrad 0.97 0.71 0.84 0.39 0.92 0.79 0.91 0.94 0.73 0.82 0.84 0.81
PromptSRC 0.97 0.84 0.86 0.43 0.91 0.78 0.95 0.94 0.74 0.83 0.87 0.83
(b) Unseen
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 0.96 0.53 0.66 0.47 0.92 0.74 0.83 0.95 0.76 0.78 0.78 0.76
CoCoOp 0.95 0.63 0.67 0.41 0.90 0.71 0.75 0.94 0.72 0.76 0.73 0.74
KgCoOp 0.96 0.56 0.69 0.40 0.92 0.74 0.78 0.95 0.72 0.75 0.75 0.75
MaPLe 0.95 0.60 0.67 0.33 0.91 0.71 0.75 0.95 0.72 0.78 0.78 0.74
ProGrad 0.95 0.51 0.66 0.38 0.88 0.68 0.77 0.94 0.69 0.71 0.70 0.72
PromptSRC 0.95 0.65 0.69 0.35 0.91 0.72 0.83 0.95 0.74 0.78 0.79 0.76
C.4 Detailed results of main experiment
In this section, We present the detailed results on the calibration of unseen classes to verify that CAC can open-vocabulary calibration in existing prompt tuning. For comprehensive evaluation, 4 standard metrics are used in our evaluation of open-vocabulary confidence calibration: Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019) and Proximity Informed Expected Calibration Error (Xiong et al., 2023). ALl of calibration error is given by ×$\times 10^{-2}$.
Table 14: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
(a) W/O Calibration
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.59 9.45 9.18 6.57 1.79 2.10 4.94 3.42 3.22 3.55 5.28 4.64
CoCoOp 2.44 13.27 10.20 13.43 1.91 1.62 7.16 2.31 2.09 1.73 3.73 5.44
KgCoOp 1.47 5.10 7.18 10.74 1.94 1.91 4.80 3.28 3.22 1.21 2.97 3.98
MaPLe 2.45 20.77 17.18 16.52 1.16 2.99 14.67 2.59 3.01 2.10 2.41 7.80
ProGrad 2.33 5.81 16.81 10.19 1.52 2.55 5.30 2.89 3.46 1.06 3.48 5.04
PromptSRC 1.85 4.83 7.79 15.41 1.53 1.63 5.92 2.98 1.87 0.94 2.47 4.29
(b) DAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.59 9.45 9.18 6.57 1.79 2.10 4.94 3.42 3.22 3.55 5.28 4.64
CoCoOp 2.18 6.26 9.30 3.52 3.81 7.85 5.94 3.88 6.42 8.28 5.26 5.70
KgCoOp 1.44 4.03 9.11 9.32 2.55 1.90 4.38 3.55 3.57 1.58 3.77 4.11
MaPLe 1.91 6.25 10.65 8.53 3.38 2.57 8.60 4.23 4.91 6.55 7.46 5.91
ProGrad 2.88 10.67 8.44 3.03 4.24 3.12 4.42 5.28 6.04 8.42 10.93 6.13
PromptSRC 2.10 4.53 9.59 12.48 2.58 1.59 5.59 3.55 2.51 2.81 2.71 4.55
(c) CAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 2.17 8.84 7.08 10.99 0.77 5.92 5.89 1.87 1.35 2.38 2.35 4.51
CoCoOp 2.45 7.80 9.71 8.62 1.74 1.54 4.21 3.48 2.68 1.20 3.18 4.24
KgCoOp 2.02 5.29 6.39 8.96 0.30 4.85 5.10 2.34 2.90 1.14 3.01 3.85
MaPLe 2.26 10.31 16.22 8.63 1.21 2.30 5.57 2.95 4.52 1.96 2.94 5.35
ProGrad 2.28 5.17 10.70 5.60 1.64 2.15 3.77 3.69 3.39 2.03 4.01 4.04
PromptSRC 1.79 4.14 7.82 7.89 0.92 2.22 5.83 2.29 1.86 1.24 2.19 3.47
Table 15: Adaptive Calibration Error (ACE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
(a) W/O Calibration
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.56 9.17 9.27 6.02 1.79 2.12 4.78 3.41 3.38 3.68 5.57 4.61
CoCoOp 2.04 13.10 10.13 13.36 1.78 1.65 7.18 2.17 1.90 1.69 3.86 5.35
KgCoOp 1.21 4.97 7.02 10.64 1.95 1.93 5.14 3.21 3.12 1.32 2.72 3.93
MaPLe 2.12 20.77 17.39 16.49 1.14 2.98 14.70 2.43 2.98 2.01 2.47 7.77
ProGrad 1.95 5.54 16.73 10.14 1.41 2.60 5.65 2.70 3.39 1.02 3.28 4.95
PromptSRC 1.54 4.89 7.69 15.36 1.48 1.72 5.85 2.81 1.88 0.99 2.39 4.24
(b) DAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.56 9.17 9.27 6.02 1.79 2.12 4.78 3.41 3.38 3.68 5.57 4.61
CoCoOp 1.43 5.96 9.18 3.84 3.80 7.85 5.96 3.72 6.32 8.28 5.22 5.60
KgCoOp 1.21 4.09 9.30 9.21 2.56 1.95 4.50 3.48 3.41 1.67 3.60 4.09
MaPLe 1.74 6.31 10.98 8.51 3.39 2.63 8.59 4.09 4.98 6.55 7.44 5.93
ProGrad 2.21 10.49 8.59 4.52 4.23 3.18 4.43 5.11 5.94 8.42 10.86 6.18
PromptSRC 1.40 4.33 9.67 12.37 2.57 1.66 5.34 3.36 2.45 2.85 2.56 4.41
(c) CAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 2.03 8.87 6.87 10.86 0.74 5.92 6.83 1.83 1.64 2.29 1.60 4.50
CoCoOp 1.82 7.92 9.6 8.43 1.65 1.62 4.9 3.3 2.68 1.2 3.27 4.22
KgCoOp 1.57 5.04 6.61 9.06 0.34 4.75 5.47 2.23 2.69 1.03 2.82 3.78
MaPLe 2.02 10.63 16.09 8.54 1.21 2.32 5.62 2.8 4.4 2.07 2.63 5.30
ProGrad 1.9 5.55 10.64 5.52 1.59 2.18 4.33 3.45 3.16 2.15 4.08 4.05
PromptSRC 1.66 5.04 7.42 7.7 0.84 2.18 5.7 2.08 1.67 1.32 1.83 3.40
Table 16: Maximum Calibration Error (MCE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
(a) W/O Calibration
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 0.49 2.84 2.45 1.75 0.62 0.49 1.03 1.20 0.73 0.68 1.11 1.22
CoCoOp 1.01 2.88 3.10 2.81 0.59 0.43 1.52 0.82 0.51 0.51 0.97 1.38
KgCoOp 0.46 1.35 2.11 2.71 0.65 0.39 1.14 1.24 0.71 0.32 0.77 1.08
MaPLe 0.84 3.71 5.59 4.07 0.46 0.83 4.33 1.04 0.73 0.65 0.66 2.08
ProGrad 1.31 1.77 4.32 2.67 0.55 0.80 1.25 1.07 0.94 0.33 1.12 1.47
PromptSRC 0.66 1.11 2.29 3.56 0.56 0.56 1.42 1.00 0.52 0.29 0.78 1.16
(b) DAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 0.49 2.84 2.45 1.75 0.62 0.49 1.03 1.20 0.73 0.68 1.11 1.22
CoCoOp 0.87 1.68 3.34 0.79 1.09 1.17 1.35 1.25 1.39 1.30 1.15 1.40
KgCoOp 0.42 1.17 3.07 2.49 0.83 0.40 1.03 1.30 0.81 0.37 1.05 1.18
MaPLe 0.62 2.01 4.60 1.83 1.07 0.51 1.53 1.45 1.27 1.11 1.82 1.62
ProGrad 1.22 2.83 2.50 0.85 1.08 0.63 1.10 1.65 1.25 1.40 2.37 1.53
PromptSRC 0.60 1.15 2.75 2.89 0.87 0.40 1.16 1.19 0.56 0.58 0.76 1.17
(c) CAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 1.02 2.39 2.40 3.06 0.21 1.80 1.17 0.60 0.30 0.83 0.52 1.30
CoCoOp 0.91 1.73 3.17 2.39 0.55 0.41 0.96 1.28 0.67 0.32 0.78 1.20
KgCoOp 0.75 1.45 2.17 2.33 0.14 1.50 1.13 0.93 0.67 0.3 0.76 1.10
MaPLe 0.74 2.49 6.22 2.47 0.48 0.67 1.12 1.17 1.04 0.42 0.94 1.61
ProGrad 1.20 1.55 3.29 1.76 0.59 0.63 1.00 1.35 0.66 0.45 1.21 1.24
PromptSRC 0.68 1.14 2.29 2.52 0.31 0.82 1.26 0.78 0.52 0.33 0.65 1.03
Table 17: Proximity-Informed Expected Calibration Error (PIECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
(a) W/O Calibration
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 3.69 14.22 10.14 10.24 2.24 2.98 7.99 3.89 4.76 4.39 7.98 6.59
CoCoOp 4.28 16.24 13.04 15.58 2.26 2.59 9.13 3.08 4.65 3.25 6.78 7.35
KgCoOp 3.85 12.05 9.75 13.32 2.40 2.78 8.15 3.86 5.21 2.90 6.69 6.45
MaPLe 4.16 22.34 20.80 18.38 1.87 3.61 15.27 3.39 4.99 3.49 6.54 9.53
ProGrad 5.01 11.67 17.25 13.17 2.26 3.30 8.86 3.63 5.44 2.98 7.98 7.41
PromptSRC 3.55 12.94 10.80 16.51 2.08 2.72 8.21 3.83 4.17 2.74 6.13 6.70
(b) DAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 3.69 14.22 10.14 10.24 2.24 2.98 7.99 3.89 4.76 4.39 7.98 6.59
CoCoOp 4.81 12.37 12.34 8.68 4.06 7.86 9.56 4.43 7.73 8.40 8.43 8.06
KgCoOp 3.88 11.51 11.52 12.26 2.87 2.75 8.19 4.11 5.39 3.05 7.31 6.62
MaPLe 4.16 11.99 16.60 11.74 3.58 3.06 10.83 4.90 6.66 6.75 9.78 8.19
ProGrad 5.43 13.87 11.16 8.70 4.50 3.67 8.60 5.83 7.26 8.54 12.65 8.20
PromptSRC 3.58 12.46 12.22 14.20 2.91 2.57 7.95 4.29 4.78 3.78 6.24 6.82
(c) CAC
Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
ZeroshotCLIP 3.59 14.01 8.49 13.39 1.33 6.26 8.15 2.77 4.07 3.43 6.16 6.51
CoCoOp 4.32 13.56 12.6 12.56 2.15 2.66 7.81 4.13 5.05 2.86 7.4 6.83
KgCoOp 3.94 12.24 9.47 11.94 1.34 5.21 8.27 3.04 5.26 2.89 6.74 6.39
MaPLe 4.04 14.86 20.58 11.16 1.93 3.06 9.18 3.74 5.96 3.23 6.8 7.69
ProGrad 4.94 10.98 12.35 10.74 2.31 3.07 8.31 4.38 5.46 3.29 8.37 6.75
PromptSRC 3.62 12.92 10.75 10.16 1.66 3.16 8.19 3.33 4.38 2.84 6.31 6.12

### v3

@@
 Vision-language models, such as CLIP (Radford et al., 2021), pre-trained on vast web-scale text-image datasets, have demonstrated impressive zero-shot capabilities and image-text alignment in various downstream image classification tasks (Deng et al., 2009; Helber et al., 2019).
-Various prompt learning methods have been proposed to enhance VLM performance on specific tasks using limited labeled data (Zhou et al., 2022b; Khattak et al., 2023a, b).
+Various prompt learning methods have been proposed to enhance VLM performance on specific tasks with a small amount of labeled data (Zhou et al., 2022b; Khattak et al., 2023a, b).
 Given CLIP’s strong zero-shot adaptability, the open-vocabulary setting has become a standard for evaluating the performance of fine-tuned VLMs, where prompts are trained on a subset of classes and evaluated on both train and unseen classes (Lee et al., 2023; Tan et al., 2024).
@@
 Confidence Calibration.
-Confidence calibration has been widely studied to ensure that the confidence levels output by models accurately reflect their true performance.
-A common approach to achieving this is by applying calibration techniques after the model has been trained.
-These calibration methods can typically be categorized into two main paradigms: scaling-based methods (Guo et al., 2017; Tomani et al., 2022; Yu et al., 2022; Xiong et al., 2023) and bin-based methods (Zadrozny & Elkan, 2001; Zhang et al., 2020; Zadrozny & Elkan, 2002; Roelofs et al., 2022).
-For scaling-based calibration, a representative method is temperature scaling (Tomani et al., 2022), which learns a single temperature value for rescaling the logits.
-ATS (Joy et al., 2023) modifies the confidence predicted by per-data-point adaptive temperature.
-With the wide use of VLMs, several works have explored calibration in VLMs (LeVine et al., 2023; Oh et al., 2023) recently.
+Confidence calibration aims to align the confidence scores predicted by models with their actual performance.
+A common strategy for achieving this is to apply calibration techniques after model training.
+These techniques can be broadly divided into two categories: scaling-based methods (Guo et al., 2017; Tomani et al., 2022; Yu et al., 2022; Xiong et al., 2023) and bin-based methods (Zadrozny & Elkan, 2001; Zhang et al., 2020; Zadrozny & Elkan, 2002; Roelofs et al., 2022).
+Among the scaling-based approaches, temperature scaling (Tomani et al., 2022) is widely used, where a single temperature parameter is learned to adjust the logits.
+ATS (Joy et al., 2023), adapts the temperature for each data point individually.
+With the growing popularity of VLMs, recent studies have also examined the calibration of these models (LeVine et al., 2023; Oh et al., 2023).
 Distance-Aware Calibration (Wang et al., 2024) estimates the scaling weights for unseen class logits based on textual features, focusing solely on changes in the textual modality.
@@
 For unseen classes calibration, we compare the SOTA method designed for CLIP, DAC (Wang et al., 2024).
-Datasets and Implementation details used in the experiments are shown in Appendix A.
+Datasets used in the experiments are shown in Appendix A.
 Implementation details.
-In the main experiments of this paper, we set $k=15$ and $\alpha=1.10$ as the default parameters and select $\lambda_{1}=0.9$ and $\lambda_{2}=1.0$ as the threshold values.
+For the main results, we use CLIP (ViT-B/16) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs.
+For the compared methods, we use their official implementations.
+In the main experiments of this paper, we set $k=15$ and $\alpha=1.10$ as the default parameters for our Contrast-Aware Weight.
+We selected $\lambda_{1}=0.9$ and $\lambda_{2}=1.0$ as the threshold values for the piecewise function.
+The rationality of these parameter choices will be validated through ablation experiments.
 Additional details about pre-trained models, hyperparameters, and implementation specifics are provided in Appendix A.
@@
 Dataset.
-Following the CoCoOp framework (Zhou et al., 2022a), we conducted evaluations of our proposed DAC along with comparison calibration methods on various image classification tasks.
+Following the CoCoOp framework (Zhou et al., 2022a), we conducted evaluations of our proposed CAC along with comparison calibration methods on various image classification tasks.
 These tasks included general object recognition using ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) datasets, fine-grained object recognition involving datasets such as Oxford Pets  (Parkhi et al., 2012), Food-101 (Bossard et al., 2014), Stanford Cars (Krause et al., 2013), Oxford Flowers102 (Nilsback & Zisserman, 2008), and FGVCAircraft (Maji et al., 2023).
@@
 Implementation details.
-We use CLIP (ViT-B/16) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs.
-We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022b).
-Following the official implementation, We list the general hyperparameters in Table 7.
-Here, we briefly introduce the corresponding exclusive hyperparameters of each VLM tuning method.
-All the methods are adopted from their official implementation.
-For CoCoOp, they do not contain other hyperparameters.
-For ProDA, we set $\lambda=0.1$ For KgCoOp, we set $\lambda=8.0$.
-For MaPLe, we set prompt depth $J$ to 0 and the language and vision prompt lengths to 2.
-For ProGrad, we set $\lambda=0.8$.
-For PromptSRC, we set deep prompting with V = T = 4.
-$\lambda_{1}=10$ and $\lambda_{2}=25$ are used in weight loss.
-For textual diversity, we use a total of $N=60$ standard prompt templates.
-We reproduce their methods in our framework based on the standard codebase.
+In our experiments, we utilize CLIP (ViT-B/16) as the pre-trained VLM and report results averaged over 3 runs.
+The model is fine-tuned in a few-shot setting (Zhou et al., 2022b) with 16 samples per class.
+The general hyperparameters are listed in Table 7, as per the official implementation.
+Below, we outline the specific hyperparameters for each VLM tuning method.
+All methods are adopted from their official implementations.
+For CoCoOp, no additional hyperparameters are required.
+In the case of ProDA, we set $\lambda=0.1$.
+For KgCoOp, $\lambda$ is set to 8.0.
+MaPLe is configured with a prompt depth $J=0$, and both the language and vision prompt lengths are set to 2.
+For ProGrad, $\lambda=0.8$ is used.
+For PromptSRC, deep prompting is set with $V=T=4$, and weight loss employs $\lambda_{1}=10$ and $\lambda_{2}=25$.
+Finally, for textual diversity, we use $N=60$ standard prompt templates.
+All methods are reproduced using the standard codebase.
 Table 7: Hyper-parameter settings for different models. All of fine-tuning method is rigorously reproduced based on the experimental details given in its paper.
@@
 Appendix B The Impact of Different backbones
-Follow DAC (Wang et al., 2024) experimental settings, we evaluated the performance of DAC and CAC on fine-tuned VLMs with different architectures using the OxfordFlowers (Nilsback & Zisserman, 2008) dataset in this section.
-Unlike accuracy, which tends to increase as the number of shots increases (Zhou et al., 2022a), calibration errors are prevalent across diverse few-shot settings.
+Following DAC (Wang et al., 2024) experimental settings, we evaluated the performance of DAC and CAC on fine-tuned VLMs with different architectures using the Oxford Flowers102 (Nilsback & Zisserman, 2008) dataset in this section.
+In contrast to accuracy, which typically exhibits an upward trend with the increase in the number of shots as reported in  (Zhou et al., 2022a), calibration errors are widespread across a variety of few-shot settings.
 As shown in Table 8, CAC demonstrates stronger cross-model generalization capabilities, achieving effective calibration across all architectures.
-Table 8: Comparison results of ECE (%) using different visual backbones on OxfordFlowers dataset. The smaller values are better.
+Table 8: Comparison results of ECE (%) using different visual backbones on Oxford Flowers102 dataset. The smaller values are better.
 Method RN50 ViT-B/32 ViT-B/16
@@
 Appendix C Detailed Experimental Results
-In this section, we illustrate the experimental results in detail.
+In this section, we provide a comprehensive details of the experimental results, with a focus on the calibration and performance of the fine-tuned VLMs.
+First, we evaluate the calibration performance on both train and unseen classes, comparing the fine-tuning results.
+Next, we further explore the calibration performance, discussing how advanced calibration techniques enhance the model’s reliability in making accurate predictions.
+Additionally, we delve into the relationship between accuracy and confidence.
+Finally, we present detailed results from the main experiments, showcasing the model’s performance across various tasks and highlighting the improvements achieved through CAC.
 C.1 The calibration performance of fine-tuned CLIP
-To showcase the miscalibration in fine-tuned VLMs, we fine-tune the pre-trained CLIP with 5 tuning methods on 11 downstream datasets and use ECE as the calibration metric.
-We report the detailed calibration results in this section.
-As is shown in Table 9(b), we can observe that distillation-based tuning (KgCoOp, ProGrad and PromptSRC) restrains the model’s confidence in train classes.
-For the calibration of the unseen class, distillation-based tuning utilizes zero-shot CLIP as the teacher model to improve the generalization ability, which generates relatively reliable predictions in unseen classes as expected.
-More importantly, we observe that fined-tuned VLMs tend to be underconfident in train classes while prone to overconfident in unseen classes.
-This phenomenon motivates us to investigate the calibration of fine-tuned VLMs further.
-Table 9: Expected Calibration Error (ECE) on diverse downstream datasets using various tuning methods for CLIP-ViT-B/16.
-The calibration performance is averaged across three variants.
-Caltech, OP, SCars, Flowers, Food, FGVC, SUN, and UCF represent Caltech101, OxfordPets, StanfordCars, Flower102, Food101, FGVCAircraft, SUN397, and UCF101.
-(a) Base
+To illustrate the miscalibration in fine-tuned VLMs, we fine-tune the pre-trained CLIP using 5 different tuning methods across 11 downstream datasets, evaluating calibration performance with ECE.
+In this section, the detailed calibration results are provided.
+As can be seen from Table 9(b), it is evident that distillation-based methods (KgCoOp, ProGrad, and PromptSRC) play a role in restricting the model’s confidence regarding train classes.
+For unseen classes, distillation-based methods leverage zero-shot CLIP as a teacher model to enhance generalization, leading to more reliable predictions for these unseen classes, as expected.
+Through comparison, we find that fine-tuned VLMs tend to be underconfident on the train classes, while they are often overconfident on unseen classes.
+This observation motivates a deeper investigation into the calibration of fine-tuned VLMs.
+Table 9: Expected Calibration Error (ECE) on diverse downstream datasets using various tuning methods for CLIP-ViT-B/16. The calibration performance is averaged across three variants.
+(a) Train
 Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
@@
 C.2 Calibrated performance for fine-tuned CLIP
-We utilize commonly used post-hoc confidence calibration for fine-tuned CLIP and expect to calibrate them well on both train and unseen classes.
-We can observe that established calibration can remedy miscalibration in train classes.
-However, such efficacy can not transfer to unseen classes.
+In our study, we apply commonly employed post-hoc confidence calibration techniques to fine-tuned VLMs, aiming to calibrate both train and unseen classes.
+Our findings reveal that the established calibration methods are capable of rectifying the miscalibration issues within the train classes.
+Nevertheless, this effectiveness fails to extend to the unseen classes.
 C.2.1 TRAIN CLASS
 Post-hoc calibration can remedy miscalibration in train classes.
-In the latest works (Zadrozny & Elkan, 2001, 2002), scaling-based methods like TS have been successfully adopted in VLM calibration, and their findings reveal that this calibration can handle miscalibration well in the close-world setting.
-As a complement, we observe that bin-based (e.g., MIR) calibration can notably reduce miscalibration in train classes.
-Post-hoc calibration on train classes can not be transferred to unseen classes.
-For instance, scaling-based methods like TS scale the logits by a single scalar temperature value.
-If fine-tuned VLMs show underconfidence in train classes, TS sharpens the logit scale before the softmax function and makes the prediction more confident.
-As a cost, the confidence distribution of fine-tuned VLMs become sharper and causes these models to display exaggerated overconfidence in unseen classes.
-for bin-based methods, they need the probabilities from train classes as the input, which are incompatible with zero-shot prediction.
+Recent studies (Zadrozny & Elkan, 2001, 2002) have successfully applied scaling-based methods like Temperature Scaling (TS) for VLM calibration, showing that such methods effectively address miscalibration in close-world environments.
+In addition, we observe that bin-based calibration methods, such as MIR, significantly reduce miscalibration in train classes.
+CAC, a calibration method designed specifically for VLMs, also generalizes well to train classes and achieves optimal calibration.
+Post-hoc calibration for train classes does not extend to unseen classes.
+For instance, scaling-based methods like TS adjust logits by a single scalar temperature value.
+When a fine-tuned VLM exhibits underconfidence in the train classes, TS sharpens the logits before the softmax function, increasing confidence in predictions.
+However, this comes at the cost of making the confidence distribution more peaked, which leads to overconfidence in unseen classes.
+Bin-based methods, on the other hand, require probabilities from train classes as input, which is incompatible with zero-probability predictions.
+This validates our motivation and highlights the rationale behind leveraging zero-shot CLIP in our approach.
 C.2.2 UNSEEN CLASS
-Table 10: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
+Table 10: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
 (a) W/O Calibration
@@
 PromptSRC 1.79 4.14 7.82 7.89 0.92 2.22 5.83 2.29 1.86 1.24 2.19 3.47
-Table 11: Expected Calibration Error (ECE) on train classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
+Table 11: Expected Calibration Error (ECE) on train classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
 (a) W/O Calibration
@@
 C.3 Accuracy & Confidence of fine-tuned CLIP
-In this section, we analyze the miscalibration from the classification performance.
-Since open-vocabulary classes are not explicitly seen during the tuning, some might believe that this miscalibration arises from rapid accuracy decreasing and leads to a larger Expected Calibration Error (ECE).
-We report the average performance on 11 datasets in Table 12(b).
-Although the classification performance on unseen classes may be comparable or lower, the average predictive confidence of fine-tuned VLMs is still relatively higher than zero-shot CLIP, which results in a larger ECE.
-The results are the same as for DAC, although fine-tuning improves the accuracy of VLM in downstream tasks, the corresponding ECE also increases.
+This section is dedicated to the analysis of miscalibration in the context of classification performance.
+Since open-vocabulary classes are not encountered during tuning, some might hypothesize that miscalibration stems from a sharp accuracy drop, leading to a higher ECE.
+We report the average performance across 11 datasets in Table 12(b).
+While classification performance on unseen classes may be comparable or lower, fine-tuned VLMs still have higher average predictive confidence than zero-shot CLIP, resulting in a larger ECE.
+This aligns with CAC results: although fine-tuning boosts accuracy, it also increases ECE.
 Table 12: Accuracy comparison of existing prompt tuning in the open-vocabulary setting.
-(a) Base
+(a) Train
 Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
@@
 Table 13: Confidence comparison of existing prompt tuning in the open-vocabulary setting.
-(a) Base
+(a) Train
 Method Caltech101 DTD EuroSAT FGVC Food101 ImageNet OF OP CARS UCF SUN AVG
@@
 C.4 Detailed results of main experiment
-In this section, We present the detailed results on the calibration of unseen classes to verify that CAC can open-vocabulary calibration in existing prompt tuning.
-For comprehensive evaluation, 4 standard metrics are used in our evaluation of open-vocabulary confidence calibration: Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019) and Proximity Informed Expected Calibration Error (Xiong et al., 2023).
-ALl of calibration error is given by ×$\times 10^{-2}$.
-Table 14: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
+This section showcases detailed results of calibration for unseen classes, aiming to illustrate that CAC facilitates open-vocabulary calibration in current prompt tuning.
+For a comprehensive evaluation, we use four standard metrics to assess open-vocabulary confidence calibration: Expected Calibration Error (ECE) (Guo et al., 2017), Maximum Calibration Error (MCE) (Guo et al., 2017), Adaptive Calibration Error (ACE) (Nixon et al., 2019), and Proximity Informed Expected Calibration Error (PIECE) (Xiong et al., 2023).
+All calibration errors are reported in $10^{-2}$.
+Table 14: Expected Calibration Error (ECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
 (a) W/O Calibration
@@
 PromptSRC 1.79 4.14 7.82 7.89 0.92 2.22 5.83 2.29 1.86 1.24 2.19 3.47
-Table 15: Adaptive Calibration Error (ACE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
+Table 15: Adaptive Calibration Error (ACE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
 (a) W/O Calibration
@@
 PromptSRC 1.66 5.04 7.42 7.7 0.84 2.18 5.7 2.08 1.67 1.32 1.83 3.40
-Table 16: Maximum Calibration Error (MCE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
+Table 16: Maximum Calibration Error (MCE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
 (a) W/O Calibration
@@
 PromptSRC 0.68 1.14 2.29 2.52 0.31 0.82 1.26 0.78 0.52 0.33 0.65 1.03
-Table 17: Proximity-Informed Expected Calibration Error (PIECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViTB/16, with calibration performance averaged across three variants.
+Table 17: Proximity-Informed Expected Calibration Error (PIECE) on unseen classes in diverse downstream datasets using various tuning methods for CLIP-ViT-B/16, with calibration performance averaged across three variants.
 (a) W/O Calibration